A camera is equipped with a lens with a focal length of 39 cm. When an object 2.6 m (260 cm) away is being photographed, what is the magnification?


1
Expert's answer
2021-11-11T09:32:05-0500

"1/a+1/b=1/f\to b=1/(1/f-1/a)=1/(1/0.39-1/2.6)=0.4588\ (m)"


"M=-b\/a=-0.4588\/2.6=-0.176" or


"M=\frac{f}{f-b}=\frac{0.39}{0.39-2.6}=-0.176" . Answer